OpenStudy (anonymous):

what is the reference angle of 5pi/4?

OpenStudy (anonymous):

The reference angle is the angle used to find the trig value when you draw the triangle in the correct quadrant. 3pi/4 and 7pi/4 and 5pi/4 all use 45 degree reference angles 11pi/3 uses a reference angle of 60 degrees. Hope this helps..

OpenStudy (anonymous):

... you could also used even multiples of \(\pi/4\) such as \(2 \pi/4=\pi/2~or~90°\), \(4 \pi/4=\pi=180°\), \(6 \pi/4=3 \pi/2=270°\) and \(8 \pi/4=2 \pi=360°\)... this will further help to locate the quadrant of \(5 \pi/4\)...

OpenStudy (noelgreco):

The reference angle is the angle to the nearest x-axis.